How many ways can you select a committee of five people from a class of 20?
Answer by super123(3)

We can select a committee of 5 pupils in a class of 20 in c(20,5)ways that is 20!/(20-5)!*5! =20*19*18*17*16
______________ 5*4*3*2*1 =15504 ways